OpenStudy (anonymous):

what number must each side of the equation 3/5x=12 be multiplied by to produce the equivalent equation x=20 -5/3 -3/5 3/5 5/3

OpenStudy (anonymous):

medal for answer

OpenStudy (anonymous):

Its D

OpenStudy (anonymous):

k can you help with a couple more?

OpenStudy (anonymous):

Multiply both sides by 5/3 5/3 * 3/5x = 5/3 * 12

OpenStudy (anonymous):

Yeah sure

OpenStudy (anonymous):

solve -1/5y=10 50 100 -100 -50

OpenStudy (anonymous):

yeah definitely -50

OpenStudy (anonymous):

solve 3lsl=18 s=-6 only s=18 or -18 s=6 only s=6 or -6

OpenStudy (anonymous):

@BlooooooBlahBLIIIIK

OpenStudy (anonymous):

idk im leaving toward C but im not totally sure on this one

OpenStudy (anonymous):

solve 2x+6-6x=30 -3 -6 3 6

OpenStudy (anonymous):

Yes 200% sure its B

OpenStudy (anonymous):

201% actually

OpenStudy (anonymous):

lol solve. 3(x-2)+5x=18 1 all real numbers 3 2

OpenStudy (anonymous):

3

OpenStudy (anonymous):

solve 7-3x=3(3-x)-2 -2 0 all real numbers 2

OpenStudy (anonymous):

well you have to subtract 7 from both sides so it would be 0

OpenStudy (anonymous):

Do you understand how i got that?

OpenStudy (anonymous):

ok that's a good thank you ^-^

OpenStudy (anonymous):

Np :) let me know if you need more help

OpenStudy (anonymous):

Can I have a metal? :)